find the exact value of the indicated term of each sequence.( show the formula and work)
4,-12,36,-108,...find a15

Answer by htmentor(1343)

find the exact value of the indicated term of each sequence.( show the formula and work)
4,-12,36,-108,...find a15
===============================
This is a geometric sequence with common ratio = -12/4 = -3
The nth term of a geometric sequence is a(n) = a*r^(n-1), where a = the 1st term
So a(15) = 4*(-3)^14 = 19131876